40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 7 10:26:36 2018, keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. photodetector_alignment.jpgdesign1.PNG
    Reply  Thu Jun 7 22:36:09 2018, not keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 00:36:54 2018, gautam, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 01:08:22 2018, johannes, Update, PSL, First light of AUX at YEND AG4395A_07-06-2018_205019.pdfPSL_AUX_SETUP.pdfAS_AUX_SETUP.pdf
       Reply  Mon Jun 18 14:32:42 2018, gautam, Update, PSL, Optics on AS table 
          Reply  Mon Jun 18 15:59:17 2018, johannes, Update, PSL, Optics on AS table 
             Reply  Mon Jun 18 16:57:54 2018, Koji, Update, PSL, Optics on AS table 
             Reply  Fri Jun 22 14:18:04 2018, Steve, Update, PSL, Optics on AS table 
Message ID: 13981     Entry time: Mon Jun 18 14:32:42 2018     In reply to: 13932     Reply to this: 13982
Author: gautam 
Type: Update 
Category: PSL 
Subject: Optics on AS table 

Yesterday, I moved the following optics:

  1. Lens in front of AS110 PD.
  2. BS splitting light between AS110 and AS55.

After moving these components around a bit, I locked them down once I was happy that the beam was pretty well centered on both of them, and also on AS110 and AS55 (measured using O'scope with single bounce from one ITM, other optics misaligned).

The beam was close to clipping on the lens mentioned in #1, probably because this wasn't checked when the 90-10 BS was installed for the AUX laser. Furthermore, I believe we are losing more than 10% of the light due to this BS. The ASDC (which is derived from AS55 PD) level is down at ~110cts as the Michelson is fringing, while it used to be ~200 cts. I will update with a power measurement shortly. But I think we should move ahead with the plan to combine the beam into the IFO's AS mode as discussed at the meeting last week.


Unrelated to this work, but c1psl and c1iscaux were keyed. 


ASDC has something weird going on with it - my main goal yesterday was to calibrate the actuators of ITMX, ITMY and BS using the Michelson. But with the Michelson locked on a dark fringe, the ASDC level changed by up to 50 counts seemingly randomly (bright fringe was ~1000 cts, I had upped the whitening gain to +21dB), even though the CCD remained clearly dark throughout. Not sure if the problem is in the readout electronics or in the PD itself.

ELOG V3.1.3-